# Active operator learning with predictive uncertainty quantification for partial differential equations

### v1

@@ Abstract
In this work, we develop a method for uncertainty quantification in deep operator networks (DeepONets) using predictive uncertainty estimates calibrated to model errors observed during training. The uncertainty framework operates using a single network, in contrast to existing ensemble approaches, and introduces minimal overhead during training and inference. We also introduce an optimized implementation for DeepONet inference (reducing evaluation times by a factor of five) to provide models well-suited for real-time applications. We evaluate the uncertainty-equipped models on a series of partial differential equation (PDE) problems, and show that the model predictions are unbiased, non-skewed, and accurately reproduce solutions to the PDEs. To assess how well the models generalize, we evaluate the network predictions and uncertainty estimates on in-distribution and out-of-distribution test datasets. We find the predictive uncertainties accurately reflect the observed model errors over a range of problems with varying complexity; simpler out-of-distribution examples are assigned low uncertainty estimates, consistent with the observed errors, while more complex out-of-distribution examples are properly assigned higher uncertainties. We also provide a statistical analysis of the predictive uncertainties and verify that these estimates are well-aligned with the observed error distributions at the tail-end of training. Finally, we demonstrate how predictive uncertainties can be used within an active learning framework to yield improvements in accuracy and data-efficiency for outer-loop optimization procedures.
@@ 1 Introduction
Understanding model uncertainty is critical for many data-driven methods; because of this, uncertainty quantification (UQ) has amassed a vast literature for quantifying what is known and what is unknown based on data availability [1, 2, 3]. When considering UQ in the context of machine learning and neural networks — which can universally approximate any continuous function [4, 5, 6] — applying traditional UQ techniques poses a challenge due to the complexity and black-box nature of the models. While accurate approximations can generally be obtained when sufficient training data is available, the resulting model predictions are still approximations, and understanding the scale and nature of the model’s errors is essential for real-world applications.
Deep operator networks (DeepONets), introduced by Lu et al. [7] in 2019, are an emerging class of neural network models that target operator approximation instead of the more conventional function approximation. Operator networks are trained on a range of distinct problems so that the trained network embodies an operator that can be used to provide rapid approximations for a variety of different problem setups. By approximating operators, these models are capable of solving entire classes of problems with a single network, whereas conventional networks are designed to solve a single problem at a time. This class of network architectures has shown significant potential with regard to predictive power, generalizability, and robustness — as seen by the wide growth and application of DeepONets in recent years [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26].
Given the prevalence of DeepONet models in practice, it is natural to ask whether such models can be enhanced to provide more information regarding the uncertainty and reliability of their predictions [21, 21, 20]. The work B-DeepONet [27] introduced a Bayesian training framework for quantifying uncertainties that result from noisy training data. This work demonstrated that the errors can be effectively quantified by incorporating probabilistic techniques into the calibration procedure for DeepONet models. While the focus of the present work does not involve noisy data, the results of B-DeepONet are promising as they show that DeepONet models can be augmented to provide accurate UQ predictions for certain problem formulations.
Yang et al. introduced UQDeepONet [28] to model uncertainty using randomized priors and leverages frequentist statistical techniques. Their work provides a scalable method for endowing DeepONets with uncertainty quantification in the presence of sufficient hardware resources, but adds significantly to the run-time costs when parallelization is not possible. This is due to the fact that the UQ framework is dependent on the evaluation of a full ensemble of independent networks; while this is generally acceptable during training, since the networks can be evaluated in parallel, it can be problematic for inference when computational resources are limited and predictions need to be made in real time.
We note that the works of Moya et al. [29, 30] were conducted in parallel with the work presented here. However, Moya’s work in 2022 focuses on applications and highlights the practical benefits of predictive uncertainty for real-time applications, demonstrating that the use of predictive uncertainty in place of ensemble Bayesian methods significantly reduces run-time costs while providing uncertainty estimates that align with Bayesian results. Moya’s 2025 work focuses on comparing conformalized DeepONets using predictive uncertainty and Bayesian ensemble uncertainty.
Pickering et al. [31] showed that DeepONet models equipped with UQ techniques can be used to improve active learning procedures. Active learning is framework for improving data efficiency of outer-loop analyses by leveraging uncertainty information to guide data acquisition. Since this can greatly reduce the costs and time spent collecting data, active learning plays an essential role in applications involving complex simulations or costly experiments [32, 33, 34, 35]. The work of Pickering employed a Monte Carlo ensemble approach for uncertainty quantification, which was shown to outperform more conventional Gaussian process surrogates.
In this work, we investigate equipping DeepONet models with predictive uncertainty capabilities analogous to those provided by ConvPDE-UQ networks [36]. We train models to approximate the solution operators for a collection of linear and non-linear partial differential equations (PDEs). We place a particular emphasis on the statistical analysis of the predictive uncertainty estimates, and we assess the uncertainty framework’s generalization capabilities on in-distribution and out-of-distribution data. While the experiments conducted in this work involve PDEs on circular and rectangular domains, the framework and network implementations can also be applied on more general domains. Finally, to demonstrate a practical use-case for the predictive uncertainty framework, we employ an uncertainty-equipped DeepONet model to accelerate an outer-loop optimization procedure using active learning. An overview of the uncertainty-equipped networks and active learning workflow is provided in Fig. 1. The main contributions of this work are summarized below.
  • Reliable uncertainty quantification in DeepONet. Our predictive uncertainty framework provides reliable uncertainty estimates calibrated to the observed model errors.
  • Efficient uncertainty estimates. The proposed framework’s predictive uncertainties provide accurate error estimates without reducing the accuracy or efficiency of DeepONet models.
  • Optimized DeepONet inference. We introduce a computational method for DeepONet inference that significantly reduces run-time costs and enables use in real-time applications.
  • UQ-based active learning for accelerated convergence. We demonstrate how predictive uncertainties can be leveraged to accelerate active learning with DeepONet surrogates. This enables active learning for problems involving function-to-function mappings, which are intractable for conventional approaches based on Gaussian process regression.
@@
Figure 1: (a) The uncertainty-equipped operator network architecture splits predictions into mean and variance estimates. We interpret network outputs as parameters for a predictive probability distribution, which are calibrated to the observed error distributions during training. (b) The uncertainty estimates provided by the operator networks are employed to help guide outer-loop active learning procedures in function spaces. The quantity of interest (QoI) and variance estimates are derived from the operator network predictions, and this information is used to guide data acquisition by balancing exploitation and exploration.
The remainder of this paper is organized as follows. In Section 2.1, we provide background for the deep operator network architecture. This is followed by a detailed summary of the proposed predictive uncertainty framework in Section 2.2. We introduce an optimized computational workflow for DeepONet inference in Section 2.3 and describe the numerical sampling procedure for generating datasets in Section 2.4. We present the experimental results for several PDE problems in Section 3 along with a detailed analysis of the predicted uncertainty estimates. We demonstrate how these predictive uncertainties can be used to accelerate active learning procedures in Section 4, and provide concluding remarks and a discussion of future research directions in Section 5.
@@ 2.1 Deep operator networks
DeepONet architectures are defined by two central components [7, 37, 38], referred to as the branch network and trunk network. The trunk network parses input coordinates corresponding to evaluation locations on the spatial (and/or temporal) domain $\Omega\subset\mathbb{R}^{d}$, and the branch network is responsible for processing information associated with the input functions, $(f,g)\in C(\Omega)\times C(\partial\Omega)$. The branch and trunk components produce vectors $b\in\mathbb{R}^{N}$ and $t\in\mathbb{R}^{N}$ with features extracted by the respective networks; these features are then combined to form the final network prediction, which is defined by the inner product $\widehat{u}(f,g)(x)=\langle\,b\,,\,t\,\rangle$ for each location $x\in\overline{\Omega}$.
@@
The component functions of the trunk network often bear a resemblance to the basis functions from finite element methods (FEM) [39, 40]. For DeepONet models, these basis functions are learned by the trunk network automatically during training, as opposed to being specified beforehand in the case of FEM calculations. The branch network assumes the role of determining the appropriate coefficients, $b$, for the family of basis functions proposed by the trunk. The resulting network approximation can then be expressed as
@@ 2.2 Equipping operator networks with predictive uncertainties
In this section, we describe how DeepONet architectures can be extended to provide predictive uncertainties following a procedure analogous to that used in the ConvPDE-UQ framework [36, 41]. Under this framework, the model provides its predictions in the form of probability distributions instead of pointwise estimates. More precisely, the network is designed to produce parameter estimates for a class of distributions (such as the mean and standard deviation for a normal distribution). These parameters are calibrated during training to capture the distribution of observed errors; the resulting model then has the ability to provide uncertainty information along with its predictions.
@@
Once an appropriate family of parameterized probability distributions is selected, the network loss is defined with respect to the associated log likelihood function, following the analogy of Gaussian process regression used in the ConvPDE-UQ framework. For example, when the network errors are observed to follow a normal distribution, the model’s proposed likelihood for the true solution is $p(u;\mu,\sigma)=(2\pi\sigma^{2})^{-1/2}\,\exp(-\tfrac{1}{2}|\mu(f,g)(x)-u(f,g)(x)|^{2}/\sigma^{2})$, and the loss function is defined by:
@@
After conducting experiments with several architecture variations (described in A.1), the network structure illustrated in Fig. 1(a) was found to provide the best performance. For the experiments presented in this work, the branch and trunk networks each consist of 5 fully-connected layers with 50 units and 30 units per layer, respectively. Each network layer uses Leaky ReLU activation functions [42], with the exception of the final layer of the branch network where no activation is applied. In order to equip the networks with predictive uncertainties, the final two layers of the branch and trunk are split to provide independent processing for the mean and uncertainty parameters. Specifically, two distinct pairs of 50 unit layers are used to produce the mean output $b$ and uncertainty output $b_{\sigma}$ of the branch network; similarly, two distinct pairs of 30 unit layers are used to produce the trunk outputs $t$ and $t_{\sigma}$. The final network outputs for the distribution parameters are then defined by the inner products: $\mu(f,g)=\langle b,t\rangle$ and $\log\sigma(f,g)=\langle b_{\sigma},t_{\sigma}\rangle$. For the experimental results presented in this work, the number of basis functions $N$ was taken to be $150$; however, the performance of the models remain essentially unchanged for any choice of $N>50$.
@@ 2.3 Precomputation to decrease computational cost
When deploying DeepONet models for real-time applications, it is important to modify the computational workflow based on the target application. We consider two common use-cases where redundant calculations can be avoided to achieve significantly faster inference speeds. We first consider the construction of light-weight surrogate models for specific PDE solutions under the assumption of fixed input functions. This is followed by an overview of how inference can be performed efficiently on the full spatial domain for arbitrary input functions.
@@ 2.3.1 Precomputing branch weights
In this setting, we note that the branch output $b$ remains constant for any choice of evaluation location $x$, since the branch depends only on the fixed input data $(f,g)$. By precomputing the branch weights, the solution $\widehat{u}(f,g)$ can be evaluated at arbitrary evaluation locations $x$ using just a single forward-pass of the trunk network. In this context, the DeepONet framework can be seen as a natural generalization of the conventional function-approximator networks [43]; i.e., once the branch weights are fixed, the DeepONet framework yields fast neural network surrogates [^2] The key practical difference is that function-approximator networks need to be re-trained whenever the input data $(f,g)$ is modified, while DeepONet surrogates can be obtained by simply recomputing the branch weights for the new inputs (requiring just one forward pass of the branch network)..

### v3

@@ Abstract
With the increased prevalence of neural operators being used to provide rapid solutions to partial differential equations (PDEs), understanding the accuracy of model predictions and the associated error levels is necessary for deploying reliable surrogate models in scientific applications. Existing uncertainty quantification (UQ) frameworks employ ensembles or Bayesian methods, which can incur substantial computational costs during both training and inference. We propose a lightweight predictive UQ method tailored for Deep operator networks (DeepONets) that also generalizes to other operator networks. Numerical experiments on linear and nonlinear PDEs demonstrate that the framework’s uncertainty estimates are unbiased and provide accurate out-of-distribution uncertainty predictions with a sufficiently large training dataset. Our framework provides fast inference and uncertainty estimates that can efficiently drive outer-loop analyses that would be prohibitively expensive with conventional solvers. We demonstrate how predictive uncertainties can be used in the context of Bayesian optimization and active learning problems to yield improvements in accuracy and data-efficiency for outer-loop optimization procedures. In the active learning setup, we extend the framework to Fourier Neural Operators (FNO) and describe a generalized method for other operator networks. To enable real-time deployment, we introduce an inference strategy based on precomputed trunk outputs and a sparse placement matrix, reducing evaluation time by more than a factor of five. Our method provides a practical route to uncertainty-aware operator learning in time-sensitive settings.
@@ 1 Introduction
Understanding and managing uncertainty is essential for data-driven models in high-consequence, real-time environments. This has motivated extensive research on uncertainty quantification (UQ), which estimates confidence in predictions based on data availability Smith [2013], Sullivan [2015], Karniadakis et al. [2021]. At the same time, scientific and engineering teams are quickly adopting neural surrogates to accelerate demanding tasks such as solving PDEs and performing adaptive simulations Zhou and Barati Farimani [2025]. These models offer speed and flexibility, but ensuring their predictions are trustworthy via traditional UQ remains a challenge due to high computational costs and the black-box nature of neural surrogates.
Neural surrogates leverage the Universal Approximation Theorem, which states that neural networks can approximate any continuous function under mild conditions Cybenko [1989], Hornik et al. [1989], Leshno et al. [1993]. Building on this, operator networks such as DeepONet Lu et al. [2021] and Fourier Neural Operators (FNO) Li et al. [2020a] extend the idea by learning solution operators for entire families of PDEs with a single trained model. When using neural surrogates, however, accurate approximations from large datasets are not enough; without trustworthy uncertainty information, surrogates can mislead downstream decisions critical for creating effective control policies and guiding experimental design. This undermines the robustness of adaptive workflows, highlighting a pressing open question: how can these models be augmented with computationally efficient and reliable uncertainty estimates to support trustworthy decision-making Zhu et al. [2023], Lee et al. [2024]?
Current UQ techniques for operator networks frequently use ensembles or Bayesian training schemes. Randomized-prior techniques like UQDeepONet Yang et al. [2022] and Bayesian DeepONet Lin et al. [2022] offer useful uncertainty estimates but have significant computational overhead, particularly during inference time when parallelism is constrained. Ensemble-based UQ evaluation costs scale linearly with ensemble size and as such, is generally impractical for real-time applications without access to considerable high-performance computing resources. As DeepONets have been promoted as fast surrogates, with work like POD-DeepONet Lu et al. [2022] targeting even further speed improvements, introducing computational bottlenecks to facilitate UQ fundamentally undermines their practical utility.
These limitations are particularly important because the uses of uncertainty quantification extends beyond error estimation; it supports adaptive procedures such as model improvement via active learning (AL) and target optimization via Bayesian optimization (BO). Pickering et al. Pickering et al. [2022] demonstrated that DeepONet models equipped with UQ can guide data acquisition for AL, reducing costs in complex simulations and experiments Settles [2009], Gal et al. [2017], Houlsby et al. [2011], Karniadakis et al. [2021]. Related strategies, such as BO, leverage UQ to balance exploration and exploitation. The literature leaves a gap for lightweight, predictive uncertainty methods that calibrate to model error in deterministic-training regimes and integrate seamlessly with outer-loop tasks.
@@
Figure 1: (a) The uncertainty-equipped operator network architecture splits predictions into mean and variance estimates. Depending on the operator network used, the uncertainty predictions could be integrated into the network like for our DeepONet, or an external (secondary) network such as for FNO. Which framework is used will depend on network architecture, but generalizes to any operator network. We interpret network outputs as parameters for a predictive probability distribution, which are calibrated to the observed error distributions during training. (b) The uncertainty estimates provided by the operator networks are employed to help guide outer-loop data aquisition procedures in function spaces. The quantity of interest (QoI) and variance estimates are derived from the operator network predictions, and this information is used to guide data acquisition by balancing exploitation and exploration.
In this work, we introduce a generalizable predictive UQ framework for operator networks, based on ConvPDE-UQ networks Winovich et al. [2019] and developed in parallel with the works of Moya et al. Moya et al. [2023, 2025]. In contrast to Moya’s 2025 work comparing conformalized DeepONets with Bayesian ensembles and other works in the literature, our framework extends predictive UQ to generalized operators for regression tasks. Additionally, we integrate our network into Bayesian-optimization and active-learning pipelines, with examples using DeepONet and FNO. We introduce a novel placement matrix for more efficient structured outputs during model deployment and provide a detailed architecture for DeepONet that delivers calibrated variance estimates using a single network. Predictive uncertainty avoids ensembles and keeps overhead low during both training and inference. The framework interprets network outputs as parameters of a predictive distribution and calibrates these parameters to observed error distributions via a negative log-likelihood (NLL) loss. To separate mean and uncertainty predictions, we split the final layers of the branch and trunk and use a log-$\sigma$ parameterization for stability.
We validate our approach on linear and nonlinear PDEs and assess generalization on both in-distribution and out-of-distribution (OoD) function spaces. The predictive uncertainty estimates are well-aligned with observed errors near the tail-end of training, remain unbiased and non-skewed, and increase appropriately in more challenging OoD regimes while staying low in simpler cases. Finally, we demonstrate how predictive UQ accelerates outer-loop procedures in a visually intuitive BO example in 1D time with two peaks in parameter space as well as a more demanding AL setup for a 2D advection problem where we compare traditional UQ methods (Bayesian networks and Ensembles) to ours. We finally showcase our method in an AL scanario where traditional UQ is too expensive, a 2D wave equation on a star-shaped domain, showing improved data efficiency and accuracy relative to baselines without UQ. An overview of the uncertainty-equipped networks and outer-loop workflows is provided in Fig. 1, with distinctions for our optimized single network DeepONet (integrated UQ) and using secondary networks for architectures like FNO (external UQ).
The main contributions of this work are summarized below.
  • Reliable predictive UQ for DeepONet. A single-network framework that calibrates uncertainty estimates to observed model errors without ensembles, yielding well-aligned uncertainty and mean predictions.
  • Generalized UQ framework for operator networks. Implementation details for how our framework can be extended to general operator learning models, with an example using FNO.
  • Efficient inference. An optimized inference method with precomputed trunk outputs and a sparse placement matrix for structured outputs.
  • Generalization to OoD data. Predictive uncertainties that reflect changing difficulty across function-space regimes, increasing on harder OoD examples and remaining low on simpler examples.
  • Accelerated outer-loop learning. Integration of predictive UQ with BO and AL to improve data-efficiency and accuracy.
The remainder of this paper is organized as follows. Section 2.1 reviews DeepONet. Section 2.2 details the predictive uncertainty framework. Section 2.3 introduces the optimized inference workflow and placement matrix. Section 2.4 describes the data generation procedure. Section 3 presents experimental results and uncertainty analysis. Sections 4 and  5.2 demonstrate BO and AL with predictive UQ, respectively. We conclude with limitations and future directions in Section 6.
@@ 2.1 Deep operator networks
DeepONet architectures are defined by two central components Lu et al. [2021], Jin et al. [2022], Lu et al. [2022], referred to as the branch network and trunk network. The trunk network parses input coordinates corresponding to evaluation locations on the spatial (and/or temporal) domain $\Omega\subset\mathbb{R}^{d}$, and the branch network is responsible for processing information associated with the input functions, $(f,g)\in C(\Omega)\times C(\partial\Omega)$. The branch and trunk components produce vectors $b\in\mathbb{R}^{N}$ and $t\in\mathbb{R}^{N}$ with features extracted by the respective networks; these features are then combined to form the final network prediction, which is defined by the inner product $\widehat{u}(f,g)(x)=\langle\,b\,,\,t\,\rangle$ for each location $x\in\overline{\Omega}$.
@@
The component functions of the trunk network often bear a resemblance to the basis functions from finite element methods (FEM) Strang and Fix [1973], Larsson and Thomée [2003]. For DeepONet models, these basis functions are learned by the trunk network automatically during training, as opposed to being specified beforehand in the case of FEM calculations. The branch network assumes the role of determining the appropriate coefficients, $b$, for the family of basis functions proposed by the trunk. The resulting network approximation can then be expressed as
@@ 2.2 Equipping operator networks with predictive uncertainties
In this section, we describe how DeepONet architectures can be extended to provide predictive uncertainties following a procedure analogous to that used in the ConvPDE-UQ framework Winovich et al. [2019], Winovich [2021]. Under this framework, the model provides its predictions in the form of probability distributions instead of pointwise estimates. More precisely, the network is designed to produce parameter estimates for a class of distributions (such as the mean and standard deviation for a normal distribution). These parameters are calibrated during training to capture the distribution of observed errors; the resulting model then has the ability to provide uncertainty information along with its predictions.
@@
Once an appropriate family of parameterized probability distributions is selected, the network loss is defined with respect to the associated log likelihood function, following the analogy of Gaussian process regression used in the ConvPDE-UQ framework Winovich et al. [2019], Winovich [2021]. For example, when the network errors are observed to follow a normal distribution, the model’s proposed likelihood for the true solution is $p(u;\mu,\sigma)=(2\pi\sigma^{2})^{-1/2}\,\exp(-\tfrac{1}{2}|\mu(f,g)(x)-u(f,g)(x)|^{2}/\sigma^{2})$, and the loss function is defined by:
@@
After conducting experiments with several architecture variations (described in A.1), the network structure illustrated in Fig. 1(a) was found to provide the best performance. For the experiments presented in this work, the branch and trunk networks each consist of 5 fully-connected layers with 50 units and 30 units per layer, respectively. Each network layer uses Leaky ReLU activation functions Maas et al. [2013], with the exception of the final layer of the branch network where no activation is applied. In order to equip the networks with predictive uncertainties, the final two layers of the branch and trunk are split to provide independent processing for the mean and uncertainty parameters. Specifically, two distinct pairs of 50 unit layers are used to produce the mean output $b$ and uncertainty output $b_{\sigma}$ of the branch network; similarly, two distinct pairs of 30 unit layers are used to produce the trunk outputs $t$ and $t_{\sigma}$. The final network outputs for the distribution parameters are then defined by the inner products: $\mu(f,g)=\langle b,t\rangle$ and $\log\sigma(f,g)=\langle b_{\sigma},t_{\sigma}\rangle$. For the experimental results presented in this work, the number of basis functions $N$ was taken to be $150$; however, the performance of the models remain essentially unchanged for any choice of $N>50$.
Although we initially developed our method for DeepONets, it is important to consider how this framework generalizes to other operator learning architectures such as FNOs Li et al. [2020a], Wen et al. [2022], Lehmann et al. [2024], Xiao et al. [2024], Grady et al. [2023], graph neural operators Li et al. [2020b], and transformer-based models Shih et al. [2025], Alkin et al. [2024], Calvello et al. [2024], Wu et al. [2024], which have seen extensive growth in the last few years Lin et al. [2021], Wang et al. [2021], Clark Di Leoni et al. [2023], Goswami et al. [2022], Mao et al. [2024]. In Sections 5.1 and 5.2, we focus on extending our framework to FNOs and leave other operator architectures for future research. Since FNOs operate in Fourier space, where evaluation grids constrain the learned mapping Lu et al. [2022], it is difficult to integrate predictive UQ estimates directly within FNO architectures. We found that introducing a secondary network for UQ, which is jointly trained with the primary model, works best. The organization of the uncertainty prediction network provides room for exploration in future research; however, this is beyond the scope of this work, and we adopt a unified training scheme for our numerical experiments.
2.2.1 Uncertainty Types and Their Role in Operator Learning
We note that traditional uncertainty quantification can be broken up into two types of uncertainty: aleatoric, which relates to inherent noise in the data, and epistemic, which reflects uncertainty due to limited knowledge or exploration of the function space. While our method primarily captures aleatoric uncertainty, we observe it can reflect epistemic-like behavior in some unexplored extrapolation regions and regions with sparse training data as seen in Section 3.4. Explicit modeling of epistemic uncertainty typically requires ensembles or Bayesian approaches, as demonstrated in recent work Guilhoto and Perdikaris [2024]. In contrast, our framework operates within the realm of aleatoric uncertainty and leverages techniques garnered from the ML community, like reinforcement learning, to help complete the “exploration” component that pure epistemic uncertainty models would handle in the BO in Section 4 and AL presented in Section 5.1 and 5.2. This strategy retains speed for practical inference while avoiding the challenge of disentangling the distinct types of uncertainty from a single deterministic model. The work of Valdenegro-Toro and Mori [2022] provides an analysis of different practical approaches for handling the disentanglement problem, but there are currently no mathematically rigorous frameworks for separating aleatoric and epistemic uncertainty, and this remains a key open problem for UQ techniques, as summarized in the conclusion of Hüllermeier and Waegeman [2021].
@@ 2.3 Precomputation to decrease computational cost
When deploying DeepONet models for real-time applications, it is important to modify the computational workflow based on the target application. We consider two common use-cases where redundant calculations can be avoided to achieve significantly faster inference speeds. We first consider the construction of light-weight surrogate models for specific PDE solutions under the assumption of fixed input functions. This is followed by an overview of how inference can be performed efficiently on the full spatial domain for arbitrary input functions, including a novel placement matrix, which, to our knowledge, has not been previously reported in literature, that enables fast and structured organization of DeepONet outputs for efficient visualization and downstream processing.
@@ 2.3.1 Precomputing branch weights
In this setting, we note that the branch output $b$ remains constant for any choice of evaluation location $x$, since the branch depends only on the fixed input data $(f,g)$. By precomputing the branch weights, the solution $\widehat{u}(f,g)$ can be evaluated at arbitrary evaluation locations $x$ using just a single forward-pass of the trunk network. In this context, the DeepONet framework can be seen as a natural generalization of the conventional function-approximator networks Chen and Chen [1995]; i.e., once the branch weights are fixed, the DeepONet framework yields fast neural network surrogates [^2] The key practical difference is that function-approximator networks need to be re-trained whenever the input data $(f,g)$ is modified, while DeepONet surrogates can be obtained by simply recomputing the branch weights for the new inputs (requiring just one forward pass of the branch network)..